• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name J. Zhang

    Log entry time 00:05:13 on June29,2010

    Entry number 328984

    keyword=Swing Shift Summary

    16:00 Shift Started with HALL in restricted acess
    17:10 Jack and smith came out from the hall
    18:45 Albert went into the hall, changed to controlled acess
    20:00 Albert finished his work. called MCC for beam permission
    20:05 Changed to empty target
    20:15 turned septum on
    21:15 Septum readback currents are all wrong.
    22:10 Called Jack for support. We decided to go down to the hall to reboot the septum magnet power supply. Called MCC for escorted acess
    22:38 Went back to beam permit
    22:53 spot++ check: Empty target, 4.3 micro A, raster 8x4 (MCC unit) gave us 4x3 mm on target.
    23:03 Move target to BeO, Atten=65, Slit=13.9, check rates
    23:32 Move to Pb#1 target
    23:35 We decided not to follow the procedure on the wiki.
    23:36 Asked MCC for 500 nA, Pb#1, rate T1=150kHz, T2=380kHz, T5=95kHz
    23:46 Changed to Ta target, 0.5 micro A
    23:56 Change to Thick Carben target